COS2A(1-1/4SIN 2 2A)=COS 6 A-SIN 6 A

COS2A(1-1/4SIN22A)=COS6A-SIN6A

Grade:11

1 Answers

Yash Baheti IIT Roorkee
9 years ago
Use these steps to solve.

Use COS 2A = Cos^2A- SIN ^2A
and on RHS, expand by using a^3-b^3= (a-b)(a^2+ab+b^2)
make wholesuquare on RHS which will become 1 and write that 1 as SIN^2 + COS^2, IN LHS wirte SIN2A as 2SIN A COS A. You will get LHS = RHS.